OpenStudy (needhelpp101):

WILL BECOME FAN && MEDAL && TESTIMONIAL!!!!!!!! Show your work. Show me how to work out he problem please A water balloon is tossed into the air with an upward velocity of 25 ft/s. Its height h(t) in ft after t seconds is given by the function h(t) = − 16t^2 + 25t + 3. 1.) After how many seconds will the balloon hit the ground? (hint: Use the quadratic formula) Answer: 1.67 sec 2.) What will the height be at t = 1 second Answer: 12 ft

OpenStudy (photon336):

okay so we know what the velocity is initially right? well I think there are two ways we could do this:

OpenStudy (needhelpp101):

ok

OpenStudy (photon336):

Finding the height at t = 1 is simple that formula h(t) => gives you the height that the object is at one second. all you need to do for the first part is just plug in t = 1 second into that equation. I don't get this. you somehow have the answer but don't know how to do it?

OpenStudy (needhelpp101):

exactly :( i don't know how but i got the answer from Mathway website

OpenStudy (needhelpp101):

@Photon336

OpenStudy (photon336):

you see this? this equation means how far the object travels in distance. t means time in seconds. so if you want to find the distance, you just plug in t =1 into the equation below. \[h(t) = -16t^{2}+25t+3\]

OpenStudy (photon336):

h(1) = -16(1)+25(1)+3 what does this give you?

OpenStudy (needhelpp101):

Look I'm just helping my daughter with her math work can you just explain please?

OpenStudy (photon336):

\[h(1) = -16+25+3 => 9+3 => 12 ft \]

OpenStudy (photon336):

well the next thing you need to do is apply the quadratic formula \[\frac{ -b~\pm~\sqrt{b^{2}-4ac} }{ 2a } => \frac{ -25~\pm~\sqrt{(25)^{2}-4(-16)(3)} }{ 2(-16) }\] all you have to do is evaluate this to get your answer \[\frac{ (-25-(28)) }{ 2(-16) } = t \]